Using a table, sketch the parametric equations, including the orientation arrows.


x = t^2 -2t

y = -t2 + 4t

0 ≤ t ≤ 3


1
Expert's answer
2021-04-12T18:09:23-0400

"x=t^2-2t\\y=-t^2+4t"

"0\leq t\leq3"


Graph of these two parametric equations are





Now upon solving the parametric equations we get,

"t^2=x+2t\\and\ \ \ t^2=4t-y"


Equating the above two equations:

"x+2t=4t-y\\t=\dfrac{x+y}{2}"


Putting the value of t in "x=t^2-2t"


"x=\dfrac{(x+y)^2}{4}-2\dfrac{(x+y)}{2}"


on solving,

"x^2+y^2-8x-4y+2xy=0"


Table:



and final locus of the parametric equation will be: